OpenStudy (anonymous):

Khalid drove 66 miles on 3 gallons of gasoline. Which is the number of miles he could drive on 12 gallons of gasoline?

OpenStudy (anonymous):

264

OpenStudy (anonymous):

66/3 = 22 miles per gallon on 12 gallons he can travel 12 * 22 = 264

OpenStudy (anonymous):

3 gal---- 66 miles 12 gal----- ? miles = 66*12/3 = 264